OpenStudy (anonymous):

log base sqrt2 Y = 3 + log base 2(Y+6)...Find Y

OpenStudy (tiffanymak1996):

3=log base 2 2^3= log base 2 8

OpenStudy (tiffanymak1996):

=log base 2 8 + log base 2 (Y+6) =log base 2 (8)(Y+6) so, Y=8Y+48 and Y= -48/7

ganeshie8 (ganeshie8):

\(\huge \log_{\sqrt{2}} \ Y = 3 + \log_2 (Y+6)\)

ganeshie8 (ganeshie8):

its like this ?

OpenStudy (anonymous):

ganeshie8 yes

ganeshie8 (ganeshie8):

ok first we need to bring logs to same base

ganeshie8 (ganeshie8):

\(\huge \log_{2^{1/2}} Y = 3 + \log_2 (Y+6)\) \(\huge \frac{\log_{2} Y}{1/2} \ = 3 + \log_2 (Y+6) \) \(\huge 2\log_2 Y - \log_2 (Y+6) = 3\) \(\huge \log_2 Y^2 - \log_2 (Y+6) = 3\)

ganeshie8 (ganeshie8):

so far makes sense ? we have used the log property : \(\log_{a^m} b = \frac{\log_a b}{m}\)

OpenStudy (anonymous):

i wasn't taught this log property. Mine working was \[\log_{\sqrt{2}}Y =\log_{2}Y \div \log_{2}\sqrt{2} \]

ganeshie8 (ganeshie8):

hmm we can use this also. but this makes the solution bit messy its good you learn that property

OpenStudy (anonymous):

can u work out the answer for me? I want to tally with mine. thanks

ganeshie8 (ganeshie8):

rest of the steps are straight forward.. il put them wait

ganeshie8 (ganeshie8):

\(\huge \frac{Y^2}{(Y+6)} = 2^3\) \(\huge Y^2 = 8Y + 48\) \(\huge Y^2 - 8Y - 48 = 0\) => Y = -4, 12 since Y cannot be negative. Y = 12

OpenStudy (anonymous):

thank you...using my method i get y = -7.06 which is wrong but I don't know the right ans so shall use yours as my correction.

ganeshie8 (ganeshie8):

you can plugin 12 and check. it satisfies... but your method also should give 12

ganeshie8 (ganeshie8):

your method is more basic... it should work as well. right now im running... maybe you msg me ur work if psble... il have a look ok :D

OpenStudy (anonymous):

I will post the pic later and u check for me when free..thank you very much,

ganeshie8 (ganeshie8):

sure np :) wolf says Y=12 is correct http://www.wolframalpha.com/input/?i=3+%2B+%28log+18%29%2Flog+2+-+%28log+12%29%2F+log+%28sqrt+2%29

OpenStudy (anonymous):